A $1,000 par bond with a 4% semi-annual coupon has 15 years to maturity trades at a yield of 6%. What would be it’s price? A. $987.44 B. $805.76 C. $1,195.99 D. $804.00
Answer #1

Coupon = (0.04 * 1000) / 2 = 20

Number of periods = 15 * 2 = 30

Rate = 6% / 2 = 3%

Price = Coupon * [1 - 1 / (1 + r)n] / r + FV / (1 + r)n

Price = 20 * [1 - 1 / (1 + 0.03)30] / 0.03 + 1000 / (1 + 0.03)30

Price = 20 * 19.60044 + 411.98676

Price = $804.00